The measures of two complementary angles are 12q-9 and 8q+14. Find the measures of the angles.
Answer by rapaljer(4671)

12q-9 + 8q+14 = 90
20q + 5 = 90
20q= 85
q = 85/20 = 4.25

The angles are 12q-9= 12(4.25) - 9 = 51-9= 42 degrees
8q + 14 = 8(4.25) + 14 = 34 + 14 = 48 degrees

Check by adding 42 + 48 = 90 degrees.

R^2 at SCC